OpenStudy (anonymous):

Let f(x)=(x+2/3)squared. find all values for the variable x for which f(x)=0.

OpenStudy (bahrom7893):

x = -2

OpenStudy (bahrom7893):

because the value of that fraction is equal to zero only when the numerator, or x plus two is equal to zero. x+2 = 0, x= -2 click on Become a fan if I helped. Thanks.

OpenStudy (anonymous):

x = -2/3

OpenStudy (bahrom7893):

JJANELLE WRONG!

OpenStudy (bahrom7893):

plug it in and see for yourself.

OpenStudy (anonymous):

\[f(x)=(x+2/3)^{2}\] is the equation to which we are referring, correct? If so, f(-2)=-16/9

OpenStudy (bahrom7893):

okay, let me prove this: \[f(-2) = ((-2+2)/3)^2 = (0/3)^2=0\]

OpenStudy (bahrom7893):

oh wait, I know what you mean!

OpenStudy (bahrom7893):

\[-2^2 + (2/3)^2\] Oh, I am sorry, I was thinking about a completely different problem

OpenStudy (bahrom7893):

You are right! But it still depends on the question, he didn't isolate (2/3) by putting it in parenthesis!

OpenStudy (bahrom7893):

Please guys, use parenthesis! And jjanelle congrats on becoming a lifesaver!

OpenStudy (anonymous):

Thanks to the order of operations parentheses aren't needed in this problem. PEMDAS - Parentheses, exponents, multiplication, division, addition, and subtraction

OpenStudy (bahrom7893):

I don't know, I always use em, especially dealing with kids on yahoo answers. I mean having them can't hurt..